OpenStudy (anonymous):

sum of the infinite series 2+(2/5)+(2/25)+(2/125)

OpenStudy (anonymous):

\[\sum_{n=0}^\infty ar^n=\frac{a}{1-r}\] (for |r|<1). Your series is \[\sum_{n=0}^\infty 2\left(\frac{1}{5}\right)^n\]